OpenStudy (anonymous):

Solve and tell whether or not it is extraneous (3x-4)^2=16

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

9x^2+24x+16=16 I thnk :)

OpenStudy (anonymous):

oh hold the phone

OpenStudy (anonymous):

if \[(3x-4)^2=16\] then either \[3x-4=4\] or \[3x-4=-4\] solve each of those

OpenStudy (anonymous):

and quadratic equations do not have "extraneous" solutions

OpenStudy (anonymous):

3x=8 x=8 or 3x=0 x=0

OpenStudy (anonymous):

8/3

OpenStudy (anonymous):

ok right, but \[3x=8\\ x=8\]is silly

OpenStudy (anonymous):

yeah \(\frac{8}{3}\)

OpenStudy (anonymous):

:) Ok. Still deciding whether or not im offended...loll :D

OpenStudy (anonymous):

so would it be non-extraneous?

OpenStudy (anonymous):

guess the honeymoon is over already :(

OpenStudy (anonymous):

yes it is "non extraneous" whatever that means

OpenStudy (anonymous):

Thanks:) Hahaha, I know! I just called off the engagement

OpenStudy (anonymous):

can i have my ring back?

OpenStudy (anonymous):

Psych! I'm selling that at the pawn shop:P